OpenStudy (taepoohwae):

Please help I have no clue! Could someone please solve this problem for me by using the reverse in foil method for factoring a trinomial? 4x+5x-15

OpenStudy (comm.dan):

Do u know how to factor trinomials, at least? @taepoohwae

OpenStudy (taepoohwae):

No I do not. I need something to go on. I'm sorry it was written wrong. 4x^2+11x+6.

OpenStudy (comm.dan):

K Thx

OpenStudy (comm.dan):

I believe this not factorable.

OpenStudy (taepoohwae):

Thanks

OpenStudy (comm.dan):

No problem, if u think I am incorrect Ask @jim_thompson5910

OpenStudy (tkhunny):

Did you try it? 4x^2+11x+6. Factors of 4 1*4 2*2 Factors of 6 1*6 2*3 There are only a few to try. Try them!

OpenStudy (anikhalder):

4x^2 +11x +6 =4x^2 +8x +3x +6 =4x(x+2) +3(x+2) (4x+3)(x+2)

OpenStudy (comm.dan):

I did it with 2x and 2x, my bad.

jimthompson5910 (jim_thompson5910):

4x^2+11x+6 4x^2+3x+8x+6 ... 3 and 8 multiply to 4*6 = 24 AND they add to 11 (middle term) (4x^2+3x)+(8x+6) x(4x+3)+(8x+6) x(4x+3)+2(4x+3) (x+2)(4x+3)

OpenStudy (comm.dan):

I see what I did wrong I needed to do 4x and x.

OpenStudy (taepoohwae):

Thanks guys! This helps a lot seeing everything in steps.